# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v1

@@ Abstract
We present a comprehensive study of axion condensed neutron stars that arise in models of an exceptionally light axion that couples to quantum chromodynamics (QCD). These axions solve the strong-charge-parity (CP) problem, but have a mass-squared lighter than that due to QCD by a factor of $\varepsilon<1$. Inside dense matter, the axion potential is altered, and much of the matter in neutron stars resides in the axion condensed phase where the strong-CP parameter $\theta=\pi$ and CP remains a good symmetry. In these regions, masses and interactions of nuclei are modified, in turn changing the equation of state (EOS), structure and phenomenology of the neutron stars. We take first steps toward the study of the EOS of neutron star matter at $\theta=\pi$ within chiral effective field theory and use relativistic mean field theory to deduce the resulting changes to nuclear matter and the neutron star low-density EOS. We derive constraints on the exceptionally light axion parameter space based on observations of the thermal relaxation of accreting neutron stars, isolated neutron star cooling, and pulsar glitches, excluding the region up to $5\times 10^{-7}\lesssim\varepsilon\lesssim 0.2$ for $m_{a}\gtrsim 2\times 10^{-9}\,{\rm eV}$. We comment on potential changes to the neutron star mass-radius relationship, and discuss the possibility of novel, nuclear-density compact objects with $\theta=\pi$ that are stabilized not by gravity but by the axion potential.
1 Introduction
The quantum chromodynamics (QCD) axion is one of the most well-motivated and minimal solutions to the Strong Charge-Parity (CP) problem axion1 ; axion2 ; axion3 ; Peccei:1977ur , the puzzle of the smallness of the neutron electric dipole moment (EDM). The neutron EDM is proportional to $\bar{\theta}$, a combination of Standard Model parameters,
@@ 1 Introduction
where $\theta$ is the QCD $\theta$ angle, while $Y_{u}$ and $Y_{d}$ are the up- and down-type Yukawa matrices. The combination is experimentally bounded to be $\bar{\theta}\lesssim 10^{-10}$ based on the non-observation of the neutron electric dipole moment Abel:2020pzs . The smallness of $\bar{\theta}$ is particularly puzzling since the CP-violating angle in the Cabibbo-Kobayashi-Maskawa (CKM) matrix has been measured to be $\mathcal{O}(1)$, in agreement with theoretical expectation.
@@
where $a$ is the axion field, $f_{a}$ the axion decay constant, $\alpha_{s}$ the strong coupling constant, $G^{\mu\nu}$ the color field strength tensor, and $\tilde{G}_{\mu\nu}\equiv\varepsilon_{\mu\nu\rho\sigma}G^{\rho\sigma}/2$ its dual. The entire dependence on $\bar{\theta}$ has been placed in the gluon term after a chiral rotation of the quarks. At low energies, the axion obtains a periodic potential from the coupling to gluons, and the CP-violating angle $\bar{\theta}$ relaxes to zero during cosmic evolution of the axion field in its potential Preskill:1982cy ; Abbott:1982af ; Dine:1982ah .
The QCD axion has inspired a diverse experimental campaign ParticleDataGroup:2024cfk ; ciaran_o_hare_2020_3932430 ; Hook:2018dlk ; Adams:2022pbo . Searches for the QCD axion and axion-like particles (ALPs) through their electron and photon couplings have seen tremendous progress in recent years. The axion-nucleon coupling—the coupling that defines the QCD axion and distinguishes it from ALPs that do not solve the strong CP problem—is much more difficult to probe in the laboratory Graham:2013gfa ; Budker:2013hfa ; Arvanitaki:2014dfa ; Arvanitaki:2021wjk ; Berlin:2022mia . The sensitivity of current ground-based experiments is rapidly improving ARIADNE:2017tdd ; Abel:2017rtm ; Chang:2017ruk ; JacksonKimball:2017elr ; Roussy:2020ily ; Aybas:2021nvn ; JEDI:2022hxa ; Schulthess:2022pbp ; Zhang:2022ewz ; Winter:2024cyp ; Fan:2024pxs , but is as yet orders of magnitude away from the “QCD axion line” on which the axion decay constant $f_{a}$ and the axion mass $m_{a}^{\rm(QCD)}$ are related by the QCD prediction axion1 ,
@@
Furthermore, these laboratory searches are at present far weaker than astrophysical constraints from Big Bang Nucleosynthesis Blum:2014vsa , the evolution of supernova 1987A Chang:2018rso , stability of the Earth, the Sun and white dwarfs Hook:2017psm ; Balkin:2022qer , LIGO-Virgo-Kagra (LVK) measurements of binary neutron star inspirals Hook:2017psm ; Huang:2018pbu ; Zhang:2021mks and isolated neutron star cooling nscooling_qcdaxion2024 .
Theoretically, there is renewed interest in the parameter space of axions which solve the strong CP problem but have an exceptionally light mass, $m_{a}<m_{a}^{\rm(QCD)}$, in light of the discrete symmetry noted in Ref. Hook:2018jle . A discrete $\mathbb{Z}_{\mathcal{N}}$ symmetry can suppress the axion mass below that predicted by Eq. (3) in a technically natural way by a factor of
@@
in the large $\mathcal{N}$ limit DiLuzio:2021pxd . This motivates the study of lighter-than-expected QCD axions—hereafter referred to as exceptionally light QCD axions—that still solve the strong CP problem (for ${\mathcal{N}}$ odd). Even within this model, we note that as a small ${\theta}$ angle is itself technically natural, the very small $\varepsilon\ll 1$ parameter space, in particular $\varepsilon\lesssim 10^{-10}$, cannot be well-motivated. As we discuss below, due to their high density, neutron stars provide a unique environment to test exceptional axions close to the QCD line: values of $\varepsilon$ close to unity, and $\mathbb{Z}_{\mathcal{N}}$ models for modest values of $\mathcal{N}$.
@@
Observationally, the above-mentioned studies of dense astrophysical objects place strong constraints on the axion-gluon coupling parameter space, and hence $\varepsilon$, especially at small axion masses. These constraints stem from the effect first noted in Ref. Hook:2017psm . The axion potential is modified in the presence of matter, which suppresses the vacuum QCD contribution. For small $\varepsilon$, there is a critical density above which a large dense object can contain a region with negative effective axion mass which, in turn, shifts the potential minimum from $\bar{\theta}=0$ to $\bar{\theta}=\pi$ inside the object. For Earth and solar core densities, the $\bar{\theta}=\pi$ region emerges for $\varepsilon\lesssim 10^{-13}$, while in white dwarfs, $\bar{\theta}=\pi$ occurs at $\varepsilon\lesssim 10^{-7}$ Hook:2017psm . Studies of the properties and stability of these objects have led to the most stringent constraint on the parameter space of $\varepsilon\lesssim 2\times 10^{-7}$ for $f_{a}\lesssim 10^{16}\,{\rm GeV}$ Balkin:2022qer .
@@
In this study, we focus on axion condensation in neutron stars for $\varepsilon\gtrsim 5\times 10^{-7}$. Earlier work has shown that for small axion masses ($m_{a}\lesssim 10^{-11}{\,\rm eV}$) the axion condensate extends far outside the radius of the star Hook:2017psm . In this scenario, the extended axion field mediates a force between neutron stars and is constrained by LVK observations of neutron star mergers Hook:2017psm ; Huang:2018pbu ; Zhang:2021mks . For a heavier axion with $m_{a}\gtrsim 10^{-8}{\,\rm eV}$, the axion field is largely confined inside the neutron star, and the effects of axion condensation on the neutron star interior need to be understood to predict the observational consequences.
@@
In this work, we provide the first dedicated study of the equation of state (EOS) for $15\times 10^{-7}\lesssim\varepsilon<$ 1 and its impact on neutron star structure. To calculate the EOS and phases present in the neutron star, we compare results from Chiral Effective Field Theory (ChiEFT) and relativistic mean field theory (RMFT). As noted in Ref. Hook:2017psm , the reduction in the effective mass of the axion in the presence of finite baryon density also leads to a reduction in the effective mass of nucleons. At fixed chemical potential, a smaller effective nucleon mass results in a larger baryon density. As a result, the transition from $\bar{\theta}=0$ to $\bar{\theta}=\pi$ introduces a discontinuity in the EOS. For different $\varepsilon$, this transition occurs at different locations in the neutron star, leading to a variety of observational effects. For $\varepsilon\gtrsim 0.1$, the neutron star will have an inner region with $\bar{\theta}=\pi$, separated by an axion field profile from the outside region where $\bar{\theta}=0$, while for smaller $\varepsilon$ the entire star has $\bar{\theta}=\pi$ except for a thin region near the surface. When $\varepsilon\lesssim 0.1$, $\bar{\theta}=\pi$ becomes energetically favored at a baryon chemical potential lower than that of iron ($\mu_{\rm Fe}/A\simeq m_{N}-8\,\mbox{MeV}$) and the axion pressure itself becomes important at low density, producing a state with zero pressure at non-negligible energy density. This may result in a star with a surface that is at a density above the neutron drip or above the crust-core transition density. In this case, the remaining crust phases will only appear in a thin region where $\theta$ changes from $\pi$ to 00.
@@
Recent observations of neutron stars (NSs) have provided reliable measurements of NS masses and constraints on their size. Radio observations of pulsars have now provided compelling evidence for massive neutron stars with masses greater than 2 M and constraints on the size of NSs from gravitational wave and x-ray observations suggest that typical NSs with mass $\simeq 1.4$ M is in the range 11-13 km. Taken together, these observations and recent ChiEFT-based calculations of the EOS of neutron-rich matter indicate that the pressure of dense matter inside neutron stars is relatively small up to $2n_{\rm sat}$, and increases rapidly at higher densities encountered in the neutron star inner core to support massive neutron stars Watts:2016uzu ; Drischler:2020fvz ; Drischler:2021kxf , where $n_{\rm sat}=0.16$ nucleons per fm3 is the nucleon density at nuclear saturation. A relatively soft EOS in the outer core and inner crust of a neutron star has important implications for neutron star structure, and as we discuss, axion condensation at $n_{B}\lesssim n_{\rm sat}$ could have dramatic effects on the structure and dynamics of the outer regions of the neutron star with observable consequences.
@@
The outline of this paper is as follows. In Sec. 2, we describe the most important features of axion condensation in a simple gas of non-interacting nucleons. In Sec. 3, we discuss the effects of axion condensation including the potential effects of nuclear interactions, with a focus on ChiEFT and RMFT models. Section 4 describes the impact that axion condensation has on the structure and properties of neutron stars within our nuclear model. In Sec. 5, we describe novel nuclear density objects stabilized by an axion potential rather than gravity. We present constraints in the $f_{a}$-$m_{a}$ parameter space for the phenomenological $\varepsilon$ model and the $Z_{\mathcal{N}}$ model based on neutron star observables in Sec. 6. Section 7 concludes this manuscript with a summary and outlook.
2 Axion condensation in a dilute gas of nucleons
@@
In this section, we summarize the main properties of axions discussed in Refs. Hook:2017psm ; Huang:2018pbu ; Hook:2019pbh ; Balkin:2020dsr ; Balkin:2021wea ; Zhang:2021mks ; Balkin:2021zfd ; Balkin:2022qer ; Balkin:2023xtr , which lead to the development of an axion profile around compact stellar objects with $\bar{\theta}=\pi$ inside these objects and $\bar{\theta}=0$ outside. We focus here on the scenario of a dilute nucleon gas and leave the impact of interactions to Sec. 3. We also review the main observational consequences for the Earth, the Sun, as well as white dwarfs and neutron stars. Henceforth, we take the quark mass matrix to be real and $\bar{\theta}=\theta$. Throughout the paper, we work with an axion with the following potential in vacuum
@@
and $\theta=a/f_{a}$. The minimum of this potential is at $\theta=0$ and the mass of the axion is [^1] The property of the potential in the $\mathbb{Z}_{\mathcal{N}}$ model is similar near $\theta=0$ but can be quite different near $\theta=\pm\pi$. See sec. 4.2 and  6.4 for more details.
@@
2.1 Baryon and meson masses
The spectrum of baryons and mesons and the vacuum energy are altered when $\theta\neq 0$ Vafa:1984xg ; Ubaldi:2008nf . The $\theta$ dependence of the pion mass, which is most important for our study, can be calculated by noting that the $\theta$ term in the QCD Lagrangian can be eliminated by a unitary rotation of the quark fields. When only up and down quarks are included, the transformation $u\rightarrow\exp{(i\gamma_{5}\phi_{u})}u$ and $d\rightarrow\exp{(i\gamma_{5}\phi_{d})}d$ where $\phi_{u}+\phi_{d}=\theta$ fully captures the effect of $\theta\neq 0$. From the relation $m_{\pi}^{2}\propto m_{q}=m_{u}+m_{d}$, the $\theta$ dependence of the mass of the neutral pion is Ubaldi:2008nf
@@ 2.1 Baryon and meson masses
At $\theta=\pi$, the net effect of this transformation is to change the sign of the up quark mass in the quark mass matrix and the neutral pion mass is $m_{\pi}(\theta=\pi)=(\sqrt{m_{d}-m_{u}}/\sqrt{m_{d}+m_{u}})m_{\pi}^{\rm phys}\approx 80\,{\rm MeV}$ for $m_{\pi}^{\rm phys}=138\,{\rm MeV}$, significantly lighter than at $\theta=0$. The smaller pion mass implies a longer range of the force and its implications for nuclear interactions will be discussed in Sec. 3.
@@
where $\sigma_{N}$ is the isoscalar nucleon sigma term and $\Delta\sigma$ is the isovector sigma term. Despite much interest and effort the isoscalar nucleon sigma term is not precisely known and the best estimates from Lattice QCD and phenomenology indicate that $\sigma_{N}\simeq 50\pm 10$ MeV Alarcon:2021dlz . The value of $\Delta\sigma$ can be inferred from the mass splitting of baryons. $\Delta\sigma$ can be derived from the mass splitting of the neutron and proton giving $\Delta\sigma=\left(m_{n}-m_{p}\right)_{\text{non}-\text{em}}/2\simeq 1\,{\rm MeV}$ Ubaldi2010deutBEtripAlph . The mass splitting of the nucleons and $\Xi$ baryon gives a somewhat larger value of $3.1$ MeV, CREWTHER1979123 . In what follows, we shall adopt $\sigma_{N}=50$ MeV and $\Delta\sigma=1$ MeV. For this choice, at $\theta=\pi$, the proton and neutron masses decrease by about 32 MeV and the neutron-proton mass difference increases by about 3.5 MeV.
2.2 Effective axion potential at finite density
The dependence of the axion potential on the temperature and density of the medium is essential to the understanding of the behavior of the axion field in the hot early universe and dense astrophysical objects. The temperature dependence of the axion mass in the early universe has been well-known to be important in obtaining a precise prediction of the range of QCD axion masses that can be the dark matter in the Universe GrillidiCortona:2015jxo . The density dependence of the axion mass, on the other hand, is much less explored. In the limit where the medium can be treated as a dilute gas of nucleons with density $n_{B}$, the axion potential has been shown to become shallower as the nucleon density increases, following the equation Hook:2017psm

### v2

@@ Abstract
We present a comprehensive study of axion condensed neutron stars that arise in models of an exceptionally light axion that couples to quantum chromodynamics (QCD). These axions solve the strong-charge-parity (CP) problem, but have a mass-squared lighter than that due to QCD by a factor of $\varepsilon<1$. Inside dense matter, the axion potential is altered, and much of the matter in neutron stars resides in the axion condensed phase where the strong-CP parameter $\theta=\pi$ and CP remains a good symmetry. In these regions, masses and interactions of nucleons are modified, in turn changing the equation of state (EOS), structure, and phenomenology of the neutron stars. We take the first steps toward the study of the EOS of neutron star matter at $\theta=\pi$ within chiral effective field theory and use relativistic mean field theory to deduce the resulting changes to nuclear matter and the neutron star low-density EOS. We derive constraints on the exceptionally light axion parameter space based on observations of the thermal relaxation of accreting neutron stars, isolated neutron star cooling, and pulsar glitches, excluding the region up to $5\times 10^{-7}\lesssim\varepsilon\lesssim 0.2$ for $m_{a}\gtrsim 2\times 10^{-9}\,{\rm eV}$. We comment on potential changes to the neutron star mass-radius relationship, and discuss the possibility of novel, nuclear-density compact objects with $\theta=\pi$ that are stabilized not by gravity but by the axion potential.
I Introduction
The quantum chromodynamics (QCD) axion is a well-motivated and minimal solution to the Strong Charge-Parity (CP) problem [1, 2, 3, 4], the puzzle of the smallness of the neutron electric dipole moment (EDM). The neutron EDM is proportional to $\bar{\theta}$, a combination of Standard Model parameters,
@@ I Introduction
where $\theta$ is the QCD $\theta$ angle, while $Y_{u}$ and $Y_{d}$ are the up- and down-type Yukawa matrices. The combination is experimentally bounded to be $\bar{\theta}\lesssim 10^{-10}$ based on the non-observation of the neutron electric dipole moment [5]. The smallness of $\bar{\theta}$ is particularly puzzling since the CP-violating angle in the Cabibbo-Kobayashi-Maskawa (CKM) matrix has been measured to be $\mathcal{O}(1)$, in agreement with theoretical expectation.
@@
where $a$ is the axion field, $f_{a}$ the axion decay constant, $\alpha_{s}$ the strong coupling constant, $G^{\mu\nu}$ the color field strength tensor, and $\tilde{G}_{\mu\nu}\equiv\varepsilon_{\mu\nu\rho\sigma}G^{\rho\sigma}/2$ its dual. The entire dependence on $\bar{\theta}$ has been placed in the gluon term after a chiral rotation of the quarks. At low energies, the axion obtains a periodic potential from the coupling to gluons, and the CP-violating angle $\bar{\theta}$ relaxes to zero during cosmic evolution of the axion field in its potential [6, 7, 8].
The QCD axion has inspired a diverse experimental campaign [9, 10, 11, 12]. Searches for the QCD axion and axion-like particles (ALPs) through their electron and photon couplings have seen tremendous progress in recent years. The axion-nucleon coupling—the coupling that defines the QCD axion and distinguishes it from ALPs that do not solve the strong CP problem—is much more difficult to probe in the laboratory [13, 14, 15, 16, 17]. The sensitivity of current ground-based experiments is rapidly improving [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], but is as yet orders of magnitude away from the “QCD axion line” on which the axion decay constant $f_{a}$ and the axion mass $m_{a}^{\rm(QCD)}$ are related by the QCD prediction [1],
@@
Furthermore, these laboratory searches are at present far weaker than astrophysical constraints from Big Bang Nucleosynthesis [29], the evolution of supernova 1987A [30, 31, 32, 33, 34, 35], stability of the Earth, the Sun and white dwarfs [36, 37], LIGO-Virgo-Kagra (LVK) measurements of binary neutron star inspirals [36, 38, 39] and isolated neutron star cooling [40]. See e.g. [41, 42], for recent reviews.
Theoretically, there is renewed interest in the parameter space of axions which solve the strong CP problem but have an exceptionally light mass, $m_{a}<m_{a}^{\rm(QCD)}$, in light of the discrete symmetry noted in Ref. [43]. A discrete $\mathbb{Z}_{\mathcal{N}}$ symmetry can suppress the axion mass below that predicted by Eq. (3) in a technically natural way by a factor of
@@
in the large $\mathcal{N}$ limit [44]. This motivates the study of lighter-than-expected QCD axions—hereafter referred to as exceptionally light QCD axions—that still solve the strong CP problem (for ${\mathcal{N}}$ odd). Even within this model, we note that as a small ${\theta}$ angle is itself technically natural, the very small $\varepsilon\ll 1$ parameter space, in particular $\varepsilon\lesssim 10^{-10}$, cannot be well-motivated. As we discuss below, due to their high density, neutron stars provide a unique environment to test exceptional axions close to the QCD line: values of $\varepsilon$ close to unity, and $\mathbb{Z}_{\mathcal{N}}$ models for modest values of $\mathcal{N}$.
@@
Observationally, the above-mentioned studies of dense astrophysical objects place strong constraints on the axion-gluon coupling parameter space, and hence $\varepsilon$, especially at small axion masses. These constraints stem from the effect first noted in Ref. [36]. The axion potential is modified in the presence of matter, which suppresses the vacuum QCD contribution. For small $\varepsilon$, there is a critical density above which a large, dense object can contain a region with negative effective axion mass which, in turn, shifts the potential minimum from $\bar{\theta}=0$ to $\bar{\theta}=\pi$ inside the object. For Earth and solar core densities, the $\bar{\theta}=\pi$ region emerges for $\varepsilon\lesssim 10^{-13}$, while in white dwarfs, $\bar{\theta}=\pi$ occurs at $\varepsilon\lesssim 10^{-7}$ [36]. Studies of the properties and stability of these objects have led to the most stringent constraint on the parameter space of $\varepsilon\lesssim 2\times 10^{-7}$ for $f_{a}\lesssim 10^{16}\,{\rm GeV}$ [37].
@@
In this study, we focus on axion condensation in neutron stars for $\varepsilon\gtrsim 5\times 10^{-7}$. Earlier work has shown that for small axion masses ($m_{a}\lesssim 10^{-11}{\,\rm eV}$) the axion condensate extends far outside the radius of the star [36]. In this scenario, the extended axion field mediates a force between neutron stars and is constrained by LVK observations of neutron star mergers [36, 38, 39]. For a heavier axion with $m_{a}\gtrsim 10^{-8}{\,\rm eV}$, the axion field is largely confined inside the neutron star, and the effects of axion condensation on the neutron star interior need to be understood to predict the observational consequences. Prior work has considered neutron stars in the axion condensed phase in the non-interacting Fermi gas approximation [45].
@@
In this work, we provide a dedicated study of the equation of state (EOS) for $15\times 10^{-7}\lesssim\varepsilon<$ 1 and its impact on neutron star structure. To calculate the EOS and phases present in the neutron star, we compare results from Chiral Effective Field Theory (ChiEFT) and relativistic mean field theory (RMFT). As noted in Ref. [36], the reduction in the effective mass of the axion in the presence of finite baryon density also leads to a reduction in the effective mass of nucleons. At fixed chemical potential, a smaller effective nucleon mass results in a larger baryon density. As a result, the transition from $\bar{\theta}=0$ to $\bar{\theta}=\pi$ introduces a discontinuity in the EOS. For different $\varepsilon$, this transition occurs at different locations in the neutron star, leading to a variety of observational effects. For $\varepsilon\gtrsim 0.1$, the neutron star will have an inner region with $\bar{\theta}=\pi$, separated by an axion field profile from the outside region where $\bar{\theta}=0$, while for smaller $\varepsilon$ the entire star has $\bar{\theta}=\pi$ except for a thin region near the surface. When $\varepsilon\lesssim 0.1$, $\bar{\theta}=\pi$ becomes energetically favored at a baryon chemical potential lower than that of iron ($\mu_{\rm Fe}/A\simeq m_{N}-8\,\mbox{MeV}$) and the axion pressure itself becomes important at low density, producing a state with zero pressure at non-negligible energy density. This may result in a star that is almost entirely at a density above the neutron drip density or above the crust-core transition density. In this case, the remaining crust phases will only appear in a thin region where $\theta$ changes from $\pi$ to 00.
@@
Recent observations of neutron stars (NSs) have resulted in reliable measurements of NS masses and constraints on their size. Radio observations of pulsars have now provided compelling evidence for neutron stars with masses greater than 2 M, while gravitational wave and x-ray observations suggest that typical NSs with mass $\simeq 1.4$ M have radii in the range 11-13 km. Taken together, these observations and recent ChiEFT-based calculations of the EOS of neutron-rich matter indicate that the pressure of dense matter inside neutron stars is relatively small up to $2n_{\rm sat}$, where $n_{\rm sat}=0.16$ nucleons per fm3 is the nucleon density at nuclear saturation, and increases rapidly at higher densities encountered in the neutron star inner core to support massive neutron stars [46, 47, 48]. A relatively soft EOS in the outer core and inner crust of a neutron star has important implications for neutron star structure. As we discuss, axion condensation at $n_{B}\lesssim n_{\rm sat}$ could have dramatic effects on the structure and dynamics of the outer regions of the neutron star with observable consequences.
@@
The outline of this paper is as follows. In Sec. II, we describe the most important features of axion condensation in a simple gas of non-interacting nucleons. In Sec. III, we discuss the effects of axion condensation including the potential effects of nuclear interactions, with a focus on ChiEFT and RMFT models. Section IV describes the impact that axion condensation has on the structure and properties of neutron stars within our nuclear model. In Sec. V, we describe novel nuclear density objects stabilized by an axion potential rather than gravity, first proposed in Refs. [37, 45] for a non-interacting Fermi gas. We present constraints in the $f_{a}$-$m_{a}$ parameter space for the phenomenological $\varepsilon$ model and the $Z_{\mathcal{N}}$ model based on neutron star observables in Sec. VI. Section VII concludes this manuscript with a summary and outlook.
II Axion condensation in a dilute gas of nucleons
@@
In this section, we summarize the main properties of axions discussed in Refs. [36, 38, 49, 50, 51, 39, 52, 37, 45], which lead to the development of an axion profile around compact stellar objects with $\bar{\theta}=\pi$ inside these objects and $\bar{\theta}=0$ outside. We focus here on the scenario of a dilute nucleon gas (see also [45]) and leave the impact of interactions to Sec. III. We also review the main observational consequences for the Earth, the Sun, white dwarfs and neutron stars. Henceforth, we take the quark mass matrix to be real and $\bar{\theta}=\theta$. Throughout the paper, we work with an axion with the following potential in vacuum
@@
and $\theta=a/f_{a}$. The minimum of this potential is at $\theta=0$ and the mass of the axion is [^1] The property of the potential in the $\mathbb{Z}_{\mathcal{N}}$ model is similar near $\theta=0$ but can be quite different near $\theta=\pm\pi$. See sec. IV.2 and  VI.4 for more details.
@@
II.1 Baryon and meson masses
The spectrum of baryons and mesons and the vacuum energy are altered when $\theta\neq 0$ [54, 55]. The $\theta$ dependence of the pion mass, which is most important for our study, can be calculated by noting that the $\theta$ term in the QCD Lagrangian can be eliminated by a unitary rotation of the quark fields. When only up and down quarks are included, the transformation $u\rightarrow\exp{(i\gamma_{5}\phi_{u})}u$ and $d\rightarrow\exp{(i\gamma_{5}\phi_{d})}d$ where $\phi_{u}+\phi_{d}=\theta$ fully captures the effect of $\theta\neq 0$. From the relation $m_{\pi}^{2}\propto m_{q}=m_{u}+m_{d}$, the $\theta$ dependence of the mass of the neutral pion is [55]
@@ II.1 Baryon and meson masses
At $\theta=\pi$, the net effect of this transformation is to change the sign of the up quark mass in the quark mass matrix and the neutral pion mass is $m_{\pi}(\theta=\pi)=(\sqrt{m_{d}-m_{u}}/\sqrt{m_{d}+m_{u}})m_{\pi}^{\rm phys}\approx 80\,{\rm MeV}$ for $m_{\pi}^{\rm phys}=138\,{\rm MeV}$, significantly lighter than at $\theta=0$. The smaller pion mass implies a longer range of the force; implications for nuclear interactions will be discussed in Sec. III.
@@
where $\sigma_{N}$ is the isoscalar nucleon sigma term and $\Delta\sigma$ is the isovector sigma term. Despite much interest and effort the isoscalar nucleon sigma term is not precisely known and the best estimates from Lattice QCD and phenomenology indicate that $\sigma_{N}\simeq 50\pm 10$ MeV [59]. The value of $\Delta\sigma$ can be inferred from the mass splitting of baryons. Determining $\Delta\sigma$ from the mass splitting of the neutron and proton gives $\Delta\sigma=\left(m_{n}-m_{p}\right)_{\text{non}-\text{em}}/2\simeq 1\,{\rm MeV}$ [60]. The mass splitting of the nucleons and $\Xi$ baryon gives a somewhat larger value of $3.1$ MeV [61]. In what follows, we shall adopt $\sigma_{N}=50$ MeV and $\Delta\sigma=1$ MeV. For this choice, at $\theta=\pi$, the proton and neutron masses decrease by about 32 MeV and the neutron-proton mass difference increases by about 3.5 MeV.
II.2 Effective axion potential at finite density
The dependence of the axion potential on the temperature and density of the medium is essential to understanding the behavior of the axion field in the hot, early universe and in dense astrophysical objects. The temperature dependence of the axion mass in the early universe is well-known to be key to obtaining a precise prediction of the range of QCD axion masses that can be the dark matter in the Universe [62]. The density dependence of the axion mass, on the other hand, is much less explored. In the limit where the medium can be treated as a dilute gas of nucleons with density $n_{B}$, the axion potential has been shown to become shallower as the nucleon density increases, following the equation [36]